10 gms of 57 Cobalt kept in a container beta decays with a half life of 270 days. The wt. of the material inside the container after 540 days will be nearly?

Grade:11

1 Answers

vikas askiitian expert
13 years ago

W/Wi = (1/2)T/t                         (t = half life  &  T is time , Wi = 10gms)

after 540 days

W/Wi = (1/2)540/270 = (1/2)2 =1/4

 W = Wi /4 = 10/4 = 2.5grams

 approve my ans if u like